(32_15P) is a beta-emitter with a decay constant of 5.6 x 10^-7 s^-1. 3.8g of (32_15P) is injected into a plant to serve as a tracer. Calculate the initial activity of the tracer.
1
Expert's answer
2021-02-24T12:51:17-0500
"A=\lambda N\\A=\lambda \frac{mN_A}{M}\\A=5.6\cdot10^{-7} \frac{3.8\cdot 6.02\cdot10^{23} }{31.97}=4\cdot10^{16} \ Bq"